OpenStudy (toxicsugar22):

hi can someone give me some elimatnition and substitution method to solve

OpenStudy (toxicsugar22):

i have a test tommorow on this so could u help me preper for it

OpenStudy (toxicsugar22):

please

OpenStudy (anonymous):

sure try this: 2x+3y=16 -4x+y=10

OpenStudy (anonymous):

and this: A red ticket costs 1 dollar. A blue ticket costs 3 dollars. Tom spent a total of 30 dollars on tickets and bought 16 total tickets. How many blue tickets and how many red tickets did he buy?

OpenStudy (anonymous):

is this the kind of problem you are talking about?

OpenStudy (toxicsugar22):

yes thanks

OpenStudy (toxicsugar22):

the first one is x=-1, y=6

OpenStudy (toxicsugar22):

is that right

OpenStudy (anonymous):

yep that's right

OpenStudy (anonymous):

did you try the word prob?

OpenStudy (toxicsugar22):

no i am

OpenStudy (toxicsugar22):

can u help me on that because i am so bad at word promblem

OpenStudy (anonymous):

sure. we know the price of a blue ticket and red ticket, and let's use b and r to denote how many blue tickets and red tickets are bought. Can you translate the problem into an equation for total money spent using the variables b and r first?

OpenStudy (anonymous):

write an expression for how much you spend if you buy b blue tickets and r red tickets and set that equal to the given total 30 dollars

OpenStudy (toxicsugar22):

this is not goin to be on my test but it is just questions can u give me those

OpenStudy (toxicsugar22):

please

OpenStudy (anonymous):

im not sure what you mean, you want me to give you the solution to the problem I wrote?

OpenStudy (anonymous):

or you want a different one?

OpenStudy (toxicsugar22):

a difefernt one like the first one

OpenStudy (anonymous):

ok I'll set up the word problem for you and you can solve that 3b+r=30 b+r=14

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

ok i got how u set it up

OpenStudy (toxicsugar22):

but shouldnt the 14 be 16

OpenStudy (toxicsugar22):

3b+r=30 b+r=14

OpenStudy (toxicsugar22):

shuldnt the 14be a 16

OpenStudy (toxicsugar22):

how did u get 14

OpenStudy (toxicsugar22):

im just a little confused

OpenStudy (toxicsugar22):

ok i got b=8, r-6

OpenStudy (toxicsugar22):

is it right

OpenStudy (anonymous):

woops i meant to say 14 total tickets. b=8 r=6 is right for 14 total tickets

OpenStudy (anonymous):

good you did it correctly and found a mistake haha

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

what else is on the test?

OpenStudy (toxicsugar22):

can u give me another promblem on this so im okay on this

OpenStudy (toxicsugar22):

not a word promblem

OpenStudy (anonymous):

word problem or just a system of equations?

OpenStudy (anonymous):

ok let me write one

OpenStudy (anonymous):

do you ever do more than 2 equations?

OpenStudy (anonymous):

3 equations 3 variables>?

OpenStudy (toxicsugar22):

i do u can give me that if u want

OpenStudy (anonymous):

k lets do one of those

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

2x-z+3y-18=0 4z+5x-2y=-18 -6y+z+3x+52=0

OpenStudy (anonymous):

that should be a little more difficult

OpenStudy (toxicsugar22):

u it is can u go over it step by step with me please

OpenStudy (toxicsugar22):

so i understand

OpenStudy (toxicsugar22):

iv tried this before and i have always had trouble on this

OpenStudy (anonymous):

sure start by lining up the variables: 2x+3y-z=18 5x-2y+4z=-18 3x-6y+z=-52

OpenStudy (anonymous):

now lets eliminate z first by multiplying the top equation by 2 and the bottom by -2 4x+6y-2z=18 5x-2y+4z=-18 -6x+12y-2z=104

OpenStudy (toxicsugar22):

is there an easier way to solve this do u think

OpenStudy (anonymous):

add those three equations.

OpenStudy (anonymous):

this is the typical way to do it by hand

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

is this 18 suppsosed to be 36 in the first equation, because we mulitied each one by two

OpenStudy (anonymous):

yep sorry

OpenStudy (toxicsugar22):

k

OpenStudy (anonymous):

when you add them you get one equation with x and y. so you can solve for one in terms of the other. then you plug into two of the original equations and you have a system of 2 equations and two unknowns

OpenStudy (toxicsugar22):

can u help me furthere im stuck

OpenStudy (anonymous):

ok adding all the equations you get 3x+16y=122 x=(122-16y)/3

OpenStudy (toxicsugar22):

okk i got it

OpenStudy (toxicsugar22):

thanks

OpenStudy (anonymous):

now you can plug into two of the equations and get numbers for x,y,z

OpenStudy (toxicsugar22):

i got it

OpenStudy (toxicsugar22):

can u help me with this one 5/x+5-7/x-5=12/x^2-25

OpenStudy (toxicsugar22):

i ma getting 1 as my answere

OpenStudy (anonymous):

x^2-25=(x+5)(x-5) 5*(x-5)/{(x-5)(x+5)}-7(x+5)/{(x-5)(x+5)}=12/{(x-5)(x+5)} 5x-25-7x-35=12 -2x=72 x=-36 I read the equation as: \[\frac{5}{x+5}-\frac{7}{x-5}=\frac{12}{x^2-25}\]

OpenStudy (toxicsugar22):

ok i got it it is -36

OpenStudy (toxicsugar22):

i found my mistake

OpenStudy (anonymous):

good :)

OpenStudy (toxicsugar22):

can u give me one substution mehtod promblem

OpenStudy (toxicsugar22):

please

OpenStudy (anonymous):

ok. Mary is 11 years older than her little sister. In 9 years she will be twice as old as her little sister. How old are mary and her sister?

OpenStudy (anonymous):

is this one ok?

OpenStudy (toxicsugar22):

yes hod on i will be back

OpenStudy (toxicsugar22):

i am going to take a showere and i will be back

OpenStudy (toxicsugar22):

i am going to take a showere and i will be back

OpenStudy (toxicsugar22):

what time are u at right now

OpenStudy (anonymous):

4:53

OpenStudy (anonymous):

i may not be here so the solution is Mary-13, sister-2. See if you can get that using substitution

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

thamk s

OpenStudy (toxicsugar22):

i gave u a medall did u get it

OpenStudy (anonymous):

yep i did, thanks :) good luck on the test

OpenStudy (toxicsugar22):

ok thanks